LCM,14,22,176
LCM,26,14,112
examples ofGCF,ALSO

Answer by Edwin McCravy(20086)

LCM of 14,22,176

Whether you want the LCM or the GCF, you begin by breaking
all the numbers down into a product of prime numbers:

14  = 2*7
22  = 2*11
176 = 2*88 = 2*2*44 = 2*2*2*22 = 2*2*2*2*11

 14 = 2*7
 22 = 2*11
176 = 2*2*2*2*11

To find the GCF, we color red all factors that
are common to ALL THREE

 14 = 2*7
 22 = 2*11
176 = 2*2*2*2*11

There is only one factor in common to all three,
and that is 2.  So
GCF of 14, 22, and 176 is 2 

-----------

To find the LCM, we color red what is in common 
to ANY TWO OR THREE of them:

 14 = 2*7
 22 = 2*11
176 = 2*2*2*2*11

The LCM is to form the product of ONLY ONE of each red factor
but EVERY non-red factor.

LCM = 2*11*2*2*2*2 = 1232 

------------------------------
------------------------------

LCM,26,14,112

LCM of 26,14,112 

Whether you want the LCM or the GCF, you begin by breaking
all the numbers down into a product of prime numbers:

 26 = 2*13
 14 = 2*7
112 = 2*56 = 2*2*28 = 2*2*2*14 = 2*2*2*2*7

 26 = 2*13
 14 = 2*7
112 = 2*2*2*2*7

To find the GCF, we color red all factors that
are common to ALL THREE

 26 = 2*13
 14 = 2*7
112 = 2*2*2*2*7

There is only one factor in common to all three,
and that is 2.  So
GCF of 26, 14, and 112 is 2 

-----------

To find the LCM, we color red what is in common 
to ANY TWO OR THREE of them:

 26 = 2*13
 14 = 2*7
112 = 2*2*2*2*7

The LCM is to form the product of ONLY ONE of each red factor
but EVERY non-red factor.

LCM = 2*7*13*2*2*2 = 1456 

Edwin